Crew 216 Operations Report 07Dec2019

[title Operations Report – December 7th]

[category operations-report]

Crew 216 Operations Report 07-Dec-19
SOL: 13
Name of person filing report: Mike Lawson
Non-nominal systems: NA
Generator: run
Hours run: 12
From what time last night: 1905
To what time this morning: 0700
List any additional daytime hours when the generator was run: N/A
Solar— SOC % (Before generator is run at night: 73%
Diesel Reading – 31
Station Propane Reading – 49
Ethanol Free Gasoline: N/A
Water (loft tank) (gal): 47
Water Meter (units): 14656315
Water (static tank) (gal): 450 (filled today, 2 trips to Hanksville)
Static to Loft Pump used – yes
Water in Green Hab (gal): 238
Water in Science Dome (gal): 0
Toilet tank emptied: yes
Deimos rover used: no
Hours:
Beginning charge:
Ending charge:
Currently charging:
Sojourner rover used: no
Hours:
Beginning charge:
Ending charge:
Currently charging:
Spirit rover used: no. Rover cleaned for crew turn-over.
Hours: 116.4
Beginning charge: (Before EVA): 100
Ending charge: (On return from EVA, before recharging): 100
Currently charging: yes
Opportunity rover used: no
Hours: 68.7
Beginning charge: 100
Ending charge: 100
Currently charging: yes
Curiosity rover used: no Rover cleaned for crew turn-over.
Hours: 121.4
Beginning charge: 100
Ending charge: (On return from EVA, before charging): 100%
Currently charging: yes
Notes on rovers: Opportunity and Deimos off-site for maintenance.
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
# Hours the ATVs were used today: 0
Notes on ATVs:
HabCar used and why, where? Yes, water re-supply, 2 trips to Hanksville.
CrewCar used and why, where?
General notes and comments: Simulation ended 062100 Dec 2019. Station cleaning continues. Food supply re-stocked.
Water supply re-stocked.
Summary of internet: Erratic/unstable wifi connection.
Summary of suits and radios: Suit 4 airflow is weak, possibly a battery charging issue.
Summary of Hab operations: HAB cleaning continues. Hot water line vibration is extreme when pump
engages, recommend securing it to HAB wall or risk fitting failure.
Summary of GreenHab operations: Routine watering. Greenhab cleaned for crew turn-over.
Summary of Science Dome operations: Lab cleaned for crew turn-over.
Summary of RAM operations: RAM Rover power charging stations inop, electrical issues remain unresolved. RAM cleaned for crew turn-over.
Summary of any observatory issues: Observatory cleaned for crew turn-over.
Summary of health and safety issues: NA
Questions, concerns and requests to Mission Support: NA